OpenStudy (anonymous):

Fun Land theme park has a total of 110 males and females who are employed at the park. If the number of males is 20 more than twice the number of females, how many males are employed at the park? 80 70 30 20

OpenStudy (nurali):

Females—x: x + 2x + 20 = 110 3x = 90 x = 30 Males: = 2(30) + 20 = 60 + 20 = 80